40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 23 16:41:06 2020, Koji, Update, IOO, Excess laser freq noise investigation plot.pdf
    Reply  Mon Oct 26 13:35:58 2020, Koji, Update, IOO, Excess laser freq noise investigation 6x
       Reply  Mon Oct 26 17:26:26 2020, gautam, Update, IOO, Excess laser freq noise investigation 
Message ID: 15641     Entry time: Fri Oct 23 16:41:06 2020     Reply to this: 15643
Author: Koji 
Type: Update 
Category: IOO 
Subject: Excess laser freq noise investigation 

[Koji, Rana]

We wanted to track down the excess noise seen in MC_F and other places (see the previous report by Gautam)


Setup1: The IMC was locked and MC_F signal between 500 and 1500Hz was observed. The DTT template was saved as /users/Templates/MC/MCF_noise_201023.xml

- Suspected mech resonance/jitter coupled with clipping or any other imperfections. Poked the various optics and optomechanics on the table. Basically no change. If we tap the laser chassis and the optics close to the laser source, we occasionally unlocked the IMC

- When we touched (lifted) the Innolight controller box from the shelf, for the first time we saw a significant change in the shape of the noise spectrum. The peak around the 700Hz shited towards lower frequency by a few %. Other peaks have no obvious change in the shapes and the heights.

- While observing the MC_F signal on the laptop, we went to the back of the laser controller. Placing a hand close to the fan clearly changes the peak frequency lower. By temporarily disconnecting the fan from the power supply for a short moment, the 700Hz peak could be eliminated. We also tried to see the noise level with the slow thermal servo and diagnosis DB cable disconnected, but we didn't see any significant change of the noise level.


Setup 2: Using the ALS phase tracker, we can observe the relative freq noise of the PSL laser and the ETMY AUX laser without any servo involved. This way we can freely disconnect any cables from the lasers. The measurement template for DTT was saved as /users/Templates/ALS/Y_ALS_FINE_PHASE_OUT_102320.xml

- Noise spectrum before disconnecting the cable (REF0, RMS REF1)

- The Fast PZT input to the PSL was disconnected => This made all the peaks (including the 700Hz) disappeared (REF2, RMS REF3)

- The Fast PZT input was restored as before, then the chain was disconnected at the input of the HV PZT driver (Thorlabs) => Again, this made the peaks disappeared (REF4, RMS REF5)

- The chain was disconnected at the input of the TTFSS box => Again, this made the peaks disappeared (REF6, RMS REF7)

- Disconnected the demod input and the AO cables from the IMC servo board => This made the peaks came back (REF8)

- Disconnected all the input/peripheral cables from the IMC servo board except for the connection to the TTFSS box => Still the excess noise was observed  (REF9)

- In addition to the above, the cable to the FSS box was disconnected but the ground was still touching the MC servo board => This made the peaks disappeared (REF10)

The conclusion is that the noise is injected from the main circuit of the IMC servo board.


Next time we will check if the backplane connection is doing something wrong. Also, we'll test if the presence of the RF signals does something bad to the IMC board via EMI and RFI.

We have reverted the connection and tested if we lock the IMC and Y arm. ==> We saw at least they were locked for a short period. The things are still stabilizing, but left them turned on so they keep trying to lock automatically for the night.

Attachment 1: plot.pdf  46 kB  Uploaded Fri Oct 23 18:44:57 2020  | Hide | Hide all
plot.pdf
ELOG V3.1.3-